a spherical capacitor has an inner sphere radius 10cm and an outer sphere of radius 14cm. the outer sphere is earthed and the onner sphere is given a charge of 2.5

 . the space between the concentric spheres is filled with a liquid of dielectric constant 32. determine the capacitamce of the capacitor.


1
Expert's answer
2021-04-30T11:13:32-0400

To be given in question

Inner radius(a)=10cm=0.10

meter

Outer radius (b)=14cm=0.14 meter

Charge(Q)=2.5c

"\epsilon_{r}=32"

To be asked in question

Capacitance of capacitor (C)=?

We know that ?

"C=\frac{4\pi\epsilon\epsilon_{r}(b-a)}{ab}\rightarrow(1)"

Put value in eqution (1)

"C=\frac{32\times(0.14-0.10)}{9 \times10^9
 \times 0.10\times0.14}"

"C=1.01587\times10^{-8}F"

"C=101.58\times10^{-6}F"

"C=101.58\mu F"